OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n2 - n + 2

terenzreignz (terenzreignz):

You know how to use mathematical induction? Show that it holds for n = 1 first, can you do that?

OpenStudy (anonymous):

not really

terenzreignz (terenzreignz):

Well that's a problem... first substitute 1 for n.

terenzreignz (terenzreignz):

Simplify, and then see if that number is divisible by 2.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

i got an answer of two

terenzreignz (terenzreignz):

Okay, good. Now, the second step of mathematical induction is to ASSUME that it holds for n = k Namely, we assume that k^2 - k + 2 is divisible by 2, catch me so far?

OpenStudy (anonymous):

umm ok

OpenStudy (anonymous):

how do we know its divisible

OpenStudy (anonymous):

we know its divisible for k=1, which is important, we need to see if the (k+1) fits the same format

terenzreignz (terenzreignz):

@BeastThatBeats We don't *know* that it works for n = k, we ASSUME it. Now, under that assumption, the most important step, substitute for n instead, k+1 And try to prove that it still yields an even number.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got k-k+4

OpenStudy (anonymous):

is it right??

terenzreignz (terenzreignz):

How did you get that?

OpenStudy (anonymous):

K+1^2-k+1+2 k+1-k+1+2 k-k+1+1+2 k-k+4

terenzreignz (terenzreignz):

lol no \[\Large (k+1)^2 = \color{red}?\]

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

k^2 +k+2

terenzreignz (terenzreignz):

Nope... at this level, multiplying polynomials should be second nature, tsk XD \[\Large (k+1)^2 = k^2 + \color{red}{2}k + \color{red}1 \]

OpenStudy (anonymous):

then

OpenStudy (anonymous):

umm are you there...